# Samir D. Mathur1

### v1

@@ Abstract
Electrons and positrons exist as real particles, so the vacuum must contain corresponding virtual fluctuations. Recognizing this fact leads to the process of Hawking radiation. This process allows consistency with thermodynamics, but leads to the trap of the information paradox. We escape the trap if we recognize that black holes have a large number of microstates, and thus the vacuum must have corresponding virtual fluctuations. In string theory many of these microstates have been explicitly constructed and are found to have an extended structure with no horizon (fig.2(c)). The corresponding virtual fluctuations, called vecros, give the gravitational vacuum a secret structure where it has correlations at all length scales. In this article we have given an explicit lattice model of these correlations, where the gravitational attraction of a star leads to a squeezing of these vecros; this squeezing alters the vecro distributions and converts them to on-shell fuzzball states at the black hole threshold. (In [12] it was shown that such fuzzballs would naturally acquire the Hawking temperature $T_{H}=1/8\pi GM$.) Since fuzzballs have no horizon and radiate from their surface like a normal body, we resolve the information paradox. It is crucial that vecro fluctuations are extended in size, so they are able to ‘feel around’ a region and respond to the formation of a closed trapped surface. By contrast, the traditional picture of the vacuum depicted in fig.1 sees nothing special upon horizon formation, and so cannot escape the trap of the paradox.

### v2

@@ Abstract
The picture of infall
The vecro picture tells us that the traditional black hole geometry is not an allowed solution at mass $M$; the stretching of vecros in such a geometry will always increase its energy above $M$. What we have instead at mass $M$ is a set of $Exp[S_{bek}(M)]$ fuzzball microstates with no horizon. An infalling shell of mass $M$ slows in its evolution as it reaches the horizon radius, since if it continued its collapse through the horizon then it would reach a situation where the energy in the geometry exceeded $M$. A process of tunneling starts into fuzzball states of mass $M$, and the shell turns into a fuzzball which will then radiate from its surface like a normal body.
But one can ask still ask the following questions. Is there any significance to the part of the classical black hole geometry inside the horizon? Can we change to Kruskal coordinates and follow the particles of the shell in their journey to the central singularity? In particular we can ask about the experience of an observer who was sitting at $r=0$ before the shell was sent in. Let the shell move radially inwards at the speed of light. Causality says that the observer at the origin does not feel any change to his evolution until the shell reaches $r=0$. Thus it would appear that there must be a meaning to the region $0<r<2GM$, at least for times of the order of the crossing time. What is the significance of this interior region, and how does its existence fit in with the notion that the hole is described by fuzzballs?
To understand what happens to such an observer at $r=0$, first recall the well-known phenomenon of alpha-decay. The alpha particle is trapped in a potential well in the nucleus, where it oscillates back and forth with some frequency; we can consider these oscillations as marking the ticking of a ‘clock’. Let us call this region of the potential as region A. There is a potential barrier trapping the alpha particle, but at some value of the radius $r_{0}$ the particle emerges from under the barrier and escapes to infinity. Let us call the region $r>r_{0}$ as region B.
We start with the alpha particle in region A. Thus the probability $P_{A}$ for the particle to be in A starts at the value $P_{A}=1$, while $P_{B}=0$. As time passes, $P_{A}$ decreases towards zero and $P_{B}$ increases towards unity. Let us now ask: as $P_{A}$ decreases, what does the clock ‘feel’? That is, does its period change, indicating damage to the clock? The answer is of course ‘no’, since the oscillations of the alpha particle in the well maintain their period; all that happens is that the probability of the clock existing itself goes to zero as the particle tunnels out. Thus we have a Schrodinger Cat type of situation: the alpha particle is in a superposition of two states (the states in region A and region B). We can in fact call this a Cheshire Cat situation: the alpha particle slowly vanishes from region A, and reappears in region B.
Moving closer to gravity, consider Minkowski space with an extra compact circle $S^{1}$. If there are no fermions in the theory, this metric is unstable to tunneling into a ‘bubble of nothing’ [13]. Let us place a graviton at $r=0$; this is the analogue of the ‘observer at the origin’ and the analogue of the clock in the alpha decay model. What will this graviton ‘feel’ as time evolves?
We see that the answer is exactly similar to the alpha decay case. To leading order, the overall wavefunctional of the full quantum gravity problem splits into parts. One part with probability $P_{A}(t)$ describes the graviton evolving normally in flat space in the region around $r=0$. The other part with probability $P_{B}(t)=1-P_{A}(t)$ describes a situation with no spacetime around $r=0$, and an expanding bubble with some radius $R_{b}$; the bubble metric carries small deformations corresponding to the details of the graviton wavefunction that we started with. Thus we again have a Schrodinger Cat/ Cheshire Cat situation: the part of the wavefunction described by $P_{A}$ evolves (to leading order) just as if there was no tunneling. This part of the graviton wavefunction does not get distorted in time due to the bubble nucleation; instead it is the probability $P_{A}$ of the graviton being in that state goes smoothly to zero.
A similar situation holds for the black hole. Let there be a graviton at the origin $r=0$, and let a shell of mass $M$ come in at the speed of light from infinity. Consider the situation when the shell has a radius $R=2GM+\Delta$. The wavefunctional describing the shell has a nonzero overlap with fuzzball microstates of radius $R=2GM+\Delta$. But these fuzzball microstates have a mass
$M_{fuzz}={R\over 2G}=M+{\Delta\over 2G}>M$ (6)
Thus the full quantum gravity wavefunctional is mostly concentrated around the shell state (which has a graviton inside at $r=0$). There is a small penetration ‘under the barrier’ into fuzzball states of mass $M_{fuzz}$; the details of the graviton at $r=0$ is encoded in the detailed shape of this wavefunctional.
As the shell approaches $R=2GM$, we find that $M_{fuzz}-M$ becomes smaller, and thus the part of the wavefunctional under the barrier becomes larger. At $R\approx 2GM+\lambda_{p}$, we have $M_{fuzz}=M$, and the shell can tunnel into the large family of $Exp[S_{bek}(M)]$ fuzzball microstates of mass $M$ [8] . At this stage we again have the Schrodinger Cat/Cheshire Cat picture. There is a probability $P_{A}(t)$ that we have the shell with the graviton inside, and a probability $P_{B}(t)=1-P_{A}(t)$ that there is a fuzzball state (i.e., no shell, no interior of a shell, and the details of the graviton encoded in the fuzzball wavefunctional).
With all this, we can address the puzzling aspects of the classical Penrose diagram. The Schwarzschild coordinates end outside the horizon, and cover this outside region for all times $-\infty<t<\infty$. In AdS/CFT duality, the CFT has just one time coordinate, which is analogous to the Schwarzschild time. This time captures the tunneling into fuzzballs, and so is like our system B in the above examples. But what is the significance of the Kruskal time which captures the inside of the collapsing shell? We see now that this time describes the evolution of system A, which is the part of the wavefunctional that has not yet tunneled through to fuzzballs. Thus in the fuzzball paradigm the long standing question ‘do we have both the inside and the outside of the hole?’ is resolved by the Schrodinger Cat/Cheshire Cat phenomenon, where we postulate no new physics, but find that the entropy enhanced tunneling into fuzzballs has created two branches of the wavefunctional for our macroscopic system. We need different time coordinates to describe these two branches.
Let us recall what was needed to get the above picture of the black hole interior. First, we need to have fuzzball states (i.e., microstates without horizons) to tunnel to. Second, the vacuum must have a vecro structure of fluctuations, whose deformations in the traditional black hole geometry create a state with energy $E>M$; this forces a tunneling transition from the collapsing shell to the fuzzball states. Third, we need the large Bekenstein entropy $S_{bek}$ to yield an entropy-enhancement of the tunneling so that a macroscopic system like the collapsing shell can transition to a Schrodinger Cat type state with two branches to its wavefunction. The above picture can be shown to get direct support by looking at evolution in the dual CFT; the details of this map will be presented elsewhere.
@@
Electrons and positrons exist as real particles, so the vacuum must contain corresponding virtual fluctuations. Recognizing this fact leads to the process of Hawking radiation. This process allows consistency with thermodynamics, but leads to the trap of the information paradox. We escape the trap if we recognize that black holes have a large number of microstates, and thus the vacuum must have corresponding virtual fluctuations. In string theory many of these microstates have been explicitly constructed and are found to have an extended structure with no horizon (fig.2(c)). The corresponding virtual fluctuations, called vecros, give the gravitational vacuum a secret structure where it has correlations at all length scales. In this article we have given an explicit lattice model of these correlations, where the gravitational attraction of a star leads to a squeezing of these vecros; this squeezing alters the vecro distributions and converts them to on-shell fuzzball states at the black hole threshold. (In [12] it was shown that such fuzzballs would naturally acquire the Hawking temperature $T_{H}=1/8\pi GM$.) Since fuzzballs have no horizon and radiate from their surface like a normal body, we resolve the information paradox. It is crucial that vecro fluctuations are extended in size, so they are able to ‘feel around’ a region and respond to the formation of a closed trapped surface. By contrast, the traditional picture of the vacuum depicted in fig.1 sees nothing special upon horizon formation, and so cannot escape the trap of the paradox. In the fuzzball paradigm we also find a natural explanation for the Kruskal time inside the hole as describing one part of the overall wavefunctional, while the Schwarzschild time captures the other part.